OpenStudy (anonymous):

radicals: √5/36, how to solve with fractions?

OpenStudy (dumbcow):

nothing to solve for? its already simplified

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{\frac{5}{36}}\] \[\large \frac{\sqrt{5}}{\sqrt{36}}\] \[\large \frac{\sqrt{5}}{\sqrt{6^2}}\] \[\large \frac{\sqrt{5}}{6}\] So \[\large \sqrt{\frac{5}{36}}=\frac{\sqrt{5}}{6}\]

OpenStudy (anonymous):

its that simple? thank you so much!

OpenStudy (dumbcow):

oh my bad i thought it was sqrt(5)/36

OpenStudy (dumbcow):

yeah its that easy, you can always split up a fraction into 2 radicals